# Efficiently Parameterized Neural Metriplectic Systems

### v2

@@ 3.2 Learnable metriplectic operators
Observe that the projections appearing in these expressions are the minimum necessary for guaranteeing the symmetries and degeneracy conditions necessary for metriplectic structure. In particular, conjugation by ${}P^{\perp}_{f}$ respects symmetry and ensures that both the input and output to the conjugated matrix field lie in $\mathrm{Span}(\nabla f)^{\perp}$.
@@
Note that the proposed parameterizations for ${}L,~{}M$ are not one-to-one but properly contain the set of valid nondegenerate metriplectic systems in $E,S$, since the Jacobi identity on ${}L$ necessary for a true Poisson manifold structure is not strictly enforced. For $1\leq i,j,k\leq n$, the Jacobi identity is given in components as $\sum_{\ell}L_{i\ell}\partial^{\ell}L_{jk}+L_{j\ell}\partial^{\ell}L_{ki}+L_{k\ell}\partial^{\ell}L_{ij}=0.$ However, this requirement is not often enforced in algorithms for learning general metriplectic (or even symplectic) systems, since it is considered subordinate to energy conservation and it is well known that both qualities cannot hold simultaneously in general [33].
@@ 3.3 Specific parameterizations
Now that Theorem 3.4 has provided a model class which is rich enough to express any desired metriplectic system, it remains to discuss what NMS actually learns. First, note that it is unlikely to be the case that $E,S$ are known a priori, so it is beneficial to allow these functions to be learnable alongside the governing operators ${}L,~{}M$. For simplicity, energy and entropy $E,S$ are parameterized as scalar-valued MLPs with $\tanh$ activation, although any desired architecture could be chosen for this task. The skew-symmetric matrix field ${}A=~{}A_{\mathrm{tri}}-~{}A_{\mathrm{tri}}^{\intercal}$ used to build ${}L$ is parameterized through its strictly lower-triangular part ${}A_{\mathrm{tri}}$ using a vector-valued MLP with output dimension (${n\choose 2}$, which guarantees that the mapping ${}A_{\mathrm{tri}}\mapsto~{}A$ above is bijective. Similarly, the symmetric matrix field ${}D=~{}K_{\mathrm{chol}}~{}K_{\mathrm{chol}}^{\intercal}$ is parameterized through its lower-triangular Cholesky factor ${}K_{\mathrm{chol}}$, which is a vector-valued MLP with output dimension (${r+1\choose 2}$. While this choice does not yield a bijective mapping ${}K_{\mathrm{chol}}\mapsto~{}D$ unless, e.g., ${}D$ is assumed to be positive definite with diagonal entries of fixed sign, this does not hinder the method in practice. In fact, ${}D$ should not be positive definite in general, as this is equivalent to claiming that ${}M$ is positive definite on vectors tangent to the level sets of $E$. Finally, the generic matrix field ${}B$ is parameterized as a vector-valued MLP with output dimension $nr$. Remarkably, the exterior algebraic expressions in Lemma 3.2 require less redundant operations than the corresponding matricized expressions from Theorem 3.4, and therefore the expressions from Lemma 3.2 are used when implementing NMS. Figure 3.3 summarizes this information.
@@
With these choices, the NMS parameterization of metriplectic systems requires only $(1/2)\mathopen{}\mathclose{{}\left((n+r)^{2}-(n-r)}\right)+2$ learnable scalar functions, in contrast to (${n\choose 3}+r{n\choose 2}+{r+1\choose 2}+2$ for the GNODE approach in [13] and $rn(n-1)+r^{2}+2$ for the GFINN approach in [32]. In particular, NMS is quadratic in both $n,r$ with no decrease in model expressivity, in contrast to the cubic scaling of previous methods.
@@ 3.4 Approximation and error
Leaning on Proposition 3.7 and classical universal approximation results in [15], it is further possible to establish the following error estimate also proven in Appendix A which gives an idea of the error accumulation rate that can be expected from this method.
@@ 4 Algorithm
Similar to previous approaches in [13] and [32], the learnable parameters in NMS are calibrated using data along solution trajectories to a given dynamical system. This brings up an important question regarding how much information about the system in question is realistically present in the training data. While many systems have a known metriplectic form, it is not always the case that one will know metriplectic governing equations for a given set of training data. Therefore, two approaches are considered in the experiments below corresponding to whether full or partial state information is assumed available during NMS training. More precisely, the state ${}x=(~{}x^{o},~{}x^{u})$ will be partitioned into “observable” and “unobservable” variables, where ${}x^{u}$ may be empty in the case that full state information is available. In a partially observable system ${}x^{o}$ typically contains positions and momenta while ${}x^{u}$ contains entropy or other configuration variables which are more difficult to observe during physical experiments. In both cases, NMS will learn a metriplectic system in ${}x$ according to the procedure described in Algorithm 1.
@@
Note that the batch-wise training strategy in Algorithm 1 requires initial conditions for ${}x^{u}$ in the partially observed case. There are several options for this, and two specific strategies will be considered here. Suppose the data are drawn from the training interval $[0,T]$ with initial state ${}x_{0}$ and final state ${}x_{T}$. The first strategy sets ${}x_{0}^{u}=~{}0$, ${}x_{T}^{u}=~{}1$ (where 11~{}11 is the all ones vector), and ${}x_{s}^{u}=~{}1/T$, $0<s<T$, so that the unobserved states are initially assumed to lie on a straight line. The second strategy is more sophisticated, and involves training a diffusion model to predict the distribution of ${}x^{u}$ given ${}x^{o}$. Specific details of this procedure are given in Appendix E.
@@ 5 Examples
The goal of the following experiments is to show that NMS is effective even when entropic information cannot be observed during training, yielding superior performance when compared to previous methods including GNODE, GFINN, and SPNN discussed in Section 2. The metrics considered for this purpose will be mean absolute error (MAE) and mean squared error (MSE) defined in the usual way as the average $\ell^{1}$ (resp. squared $\ell^{2}$) error between the discrete states ${}x,\tilde{~{}x}\in\mathbb{R}^{n\times n_{s}}$. For brevity, many implementation details have been omitted here and can be found in Appendix B. An additional experiment showing the effectiveness of NMS in the presence of both full and partial state information can be found in Appendix C.
@@ 5.1 Two gas containers
where $(q,p)$ are the position and momentum of the separating wall, $S_{1},S_{2}$ are the entropies of the two subsystems, and the internal energies $E_{1},E_{2}$ are determined from the Sackur-Tetrode equation for ideal gases, $S_{i}/Nk_{B}=\ln\mathopen{}\mathclose{{}\left(\hat{c}V_{i}E_{i}^{3/2}}\right),1\leq i\leq 2.$ Here, $m$ denotes the mass of the wall, $2L$ is the total length of the system, and $V_{i}$ is the volume of the $i^{\mathrm{th}}$ container. As in [13, 14] $Nk_{B}=1$ and $\alpha=0.5$ fix the characteristic macroscopic unit of entropy while $\hat{c}=102.25$ ensures the argument of the logarithm defining $E_{i}$ is dimensionless. This leads to the total entropy $S(~{}x)=S_{1}+S_{2}$ and the total energy $E(~{}x)=(1/2m)p^{2}+E_{1}(~{}x)+E_{2}(~{}x),$ which are guaranteed to be nondecreasing and constant, respectively.
@@ 5.2 Thermoelastic double pendulum
Next, consider the thermoelastic double pendulum from [23] with 10-dimensional state variable ${}x=\begin{pmatrix}~{}q_{1}&~{}q_{2}&~{}p_{1}&~{}p_{2}&S_{1}&S_{2}\end{pmatrix}^{\intercal}$, which represents a highly challenging benchmark for metriplectic methods. The equations of motion in this case are given for $1\leq i\leq 2$ as
@@
The task in this case is prediction across initial conditions. As in [32], 100 trajectories are drawn from the ranges in Appendix B and integrated over the interval $[0,40]$ with $\Delta t=0.1$, with an 80/10/10 split for training/validation/testing. Here all compared models are trained using full state information. As seen in Table 1, NMS is again the most performant, although all models struggle to approximate the dynamics over the entire training interval. It is also notable that the training time of NMS is greatly decreased relative to GNODE and GFINN due to its improved quadratic scaling; a representative study to this effect is given in Appendix D.
@@ Appendix A Proof of Theoretical Results
This will be a direct consequence of Corollary 2.2 in [15] provided we show that $\mathopen{}\mathclose{{}\left|~{}A}\right|\leq c\mathopen{}\mathclose{{}\left|~{}A}\right|_{\infty}$ for some $c>0$. To that end, if $\sigma_{1}\geq...\geq\sigma_{r}>0$ ($r\leq n$) denote the nonzero singular values of ${}A-\tilde{~{}A}$, it follows that for each ${}x\in K$,
@@
and therefore the desired inequality holds with $c=n$. Now, for any $\varepsilon>0$ it follows from [15] that there exists a two layer network $\tilde{~{}A}$ with $m$ continuous derivatives such that $\sup_{~{}x\in K}\mathopen{}\mathclose{{}\left|~{}A-\tilde{~{}A}}\right|_{\infty}<\varepsilon/n$ and $\sup_{~{}x\in K}\mathopen{}\mathclose{{}\left|\nabla^{k}~{}A-\nabla^{k}\tilde{~{}A}}\right|_{\infty}<\varepsilon/n<\varepsilon$ for all $1\leq k\leq m$. Therefore, it follows that
@@
Now, by Corollary 2.2 in [15], for any $\varepsilon>0$ there exists a two-layer neural network $\tilde{f}:K\to\mathbb{R}$ such that
@@ Appendix B Experimental and Implementation Details
This Appendix records additional details related to the numerical experiments in Section 5. For each benchmark problem, a set of trajectories is manufactured given initial conditions by simulating ODEs with known metriplectic structure. For the experiments in Table 1, only the observable variables are used to construct datasets, since entropic information is assumed to be unknown. Algorithm 2 summarizes the training of the dynamics models used for comparison with NMS.
@@
For each compared method, integrating the ODEs is done via the Dormand–Prince method (dopri5) [1] with relative tolerance $10^{-7}$ and absolute tolerance $10^{-9}$. The loss is evaluated by measuring the discrepancy between the ground truth observable states ${}x^{\text{o}}$ and the approximate observable states $\tilde{~{}x}^{\text{o}}$ in the mean absolute error (MAE) metric. The model parameters $\Theta$ (i.e., the weights and biases) are updated by using Adamax [12] with an initial learning rate of 0.01. The number of training steps is set as 30,000, and the model parameters resulting in the best performance for the validation set are chosen for testing. Specific information related to the experiments in Section 5 is given in the subsections below.
@@ B.1 Two gas containers
In the experiment, a sequence of 100,000 timesteps is generated using the Runge–Kutta 4th-order (RK4) time integrator with a step size 0.001. The initial condition is given as ${}x=(1,2,103.2874,103.2874)$ following [25]. The training/validation/test split is defined by $t_{\text{train}}=20$, $t_{\text{val}}=30$, and $t_{\text{test}}=100$. For a fair comparison, all considered models are set to have a similar number of model parameters, $\sim$2,000. The specifications of the network architectures are:

### v3

@@ 3.2 Learnable metriplectic operators
Observe that the projections appearing in these expressions are the minimum necessary for guaranteeing the symmetries and degeneracy conditions necessary for a general metriplectic structure. In particular, conjugation by ${}P^{\perp}_{f}$ respects symmetry and ensures that both the input and output to the conjugated matrix field lie in $\mathrm{Span}(\nabla f)^{\perp}$. Note that this implies that the class of nondegenerate metriplectic systems cannot be parameterized in a way that scales sub-quadratically in the state dimension $n$.
@@
As mentioned, the proposed parameterizations for ${}L,~{}M$ are not one-to-one but properly contain the set of valid nondegenerate metriplectic systems in $E,S$. This is due to both the orthogonal projections employed, which induce non-uniqueness in the ${}L,~{}M$ matrix fields, as well as the lack of a guaranteed Jacobi identity for ${}L$. Recall that the Jacobi identity is necessary for the matrix field ${}L$ to generate a true Poisson manifold structure, and hence for $\dot{~{}x}$ to be the velocity field of a truly metriplectic system. For $1\leq i,j,k\leq n$, the Jacobi identity is given in components as $\sum_{\ell}L_{i\ell}\partial^{\ell}L_{jk}+L_{j\ell}\partial^{\ell}L_{ki}+L_{k\ell}\partial^{\ell}L_{ij}=0.$ However, this requirement is not often enforced in algorithms for learning general metriplectic (or even symplectic) systems, since it is considered subordinate to energy conservation and it is well known that both qualities cannot hold simultaneously in general Zhong and Marsden (1988). Future work will compare the alternative choice of parameterizations which enforce the Jacobi identity explicitly, maintaining Poisson structure while sacrificing strong-form energy conservation.
@@ 3.3 Specific parameterizations
Now that Theorem 3.4 has provided a model class which is rich enough to express any desired metriplectic system, it remains to discuss what NMS actually learns. First, note that it is unlikely to be the case that $E,S$ are known a priori, so it is beneficial to allow these functions to be learnable alongside the governing operators ${}L,~{}M$. For simplicity, energy and entropy $E,S$ are parameterized as scalar-valued MLPs with $\tanh$ activation, although any desired architecture could be chosen for this task. The skew-symmetric matrix field ${}A=~{}A_{\mathrm{tri}}-~{}A_{\mathrm{tri}}^{\intercal}$ used to build ${}L$ is parameterized through its strictly lower-triangular part ${}A_{\mathrm{tri}}$ using a vector-valued MLP with output dimension (${n\choose 2}$, which guarantees that the mapping ${}A_{\mathrm{tri}}\mapsto~{}A$ above is bijective. Similarly, the symmetric matrix field ${}D=~{}K_{\mathrm{chol}}~{}K_{\mathrm{chol}}^{\intercal}$ is parameterized through its lower-triangular Cholesky factor ${}K_{\mathrm{chol}}$, which is a vector-valued MLP with output dimension (${r+1\choose 2}$. While this choice does not yield a bijective mapping ${}K_{\mathrm{chol}}\mapsto~{}D$ unless, e.g., ${}D$ is assumed to be positive definite with diagonal entries of fixed sign, this does not hinder the method in practice. In fact, ${}D$ should not be positive definite in general, as this is equivalent to claiming that ${}M$ is positive definite on vectors tangent to the level sets of $E$. Finally, the generic matrix field ${}B$ is parameterized as a vector-valued MLP with output dimension $nr$. Remarkably, the exterior algebraic expressions in Lemma 3.2 require less redundant operations than the corresponding matricized expressions in Theorem 3.4, and therefore the expressions from Lemma 3.2 are used when implementing NMS. This can be seen from the proof of Lemma 3.2 in Appendix A, where matricization of these expressions requires computing a term which is identically zero due to symmetry considerations. Figure 3.3 summarizes the proposed NMS architecture.
@@
With these choices, the NMS parameterization of metriplectic systems requires only $(1/2)\mathopen{}\mathclose{{}\left((n+r)^{2}-(n-r)}\right)+2$ learnable scalar functions, in contrast to (${n\choose 3}+r{n\choose 2}+{r+1\choose 2}+2$ for the GNODE approach in Lee et al. (2021) and $rn(n-1)+r^{2}+2$ for the GFINN approach in Zhang et al. (2022). In particular, NMS is quadratic in both $n,r$ with no decrease in model expressivity, in contrast to the cubic scaling of previous methods.
@@ 3.4 Approximation and error
Leaning on Proposition 3.7 and classical universal approximation results in Li (1996), it is further possible to establish the following error estimate also proven in Appendix A which gives an idea of the error accumulation rate that can be expected from this method.
@@ 4 Algorithm
Similar to previous approaches in Lee et al. (2021) and Zhang et al. (2022), the learnable parameters in NMS are calibrated using data along solution trajectories to a given dynamical system. This brings up an important question regarding how much information about the system in question is realistically present in the training data. While many systems have a known metriplectic form, it is not always the case that one will know metriplectic governing equations for a given set of training data. Therefore, two approaches are considered in the experiments below corresponding to whether full or partial state information is assumed available during NMS training. More precisely, the state ${}x=(~{}x^{o},~{}x^{u})$ will be partitioned into “observable” and “unobservable” variables, where ${}x^{u}$ may be empty in the case that full state information is available. In a partially observable system ${}x^{o}$ typically contains positions and momenta while ${}x^{u}$ contains entropy or other configuration variables which are more difficult to observe during physical experiments. In both cases, NMS will learn a metriplectic system in ${}x$ according to the procedure described in Algorithm 1.
@@
Note that the batch-wise training strategy in Algorithm 1 requires initial conditions for ${}x^{u}$ in the partially observed case. There are several options for this, and two specific strategies will be considered here. Suppose the data are drawn from the training interval $[0,T]$ with initial state ${}x_{0}$ and final state ${}x_{T}$. The first strategy sets ${}x_{0}^{u}=~{}0$, ${}x_{T}^{u}=~{}1$ (where 11~{}11 is the all ones vector), and ${}x_{s}^{u}=~{}1/s$, $0<s<T$, so that the unobserved states are initially assumed to lie on a straight line. In the absence of other information, this ensures that the entropy will be satisfy the monotonicity constraint necessary for metriplectic structure. The second strategy is more sophisticated, and involves training a diffusion model to predict the distribution of ${}x^{u}$ given ${}x^{o}$. Specific details of this procedure are given in Appendix H. Note that other strategies are also possible, such as allowing the unobserved initial states to be trainable as in Chen et al. (2020).
@@ 5 Examples
The goal of the following experiments is to show that NMS is effective even when entropic information cannot be observed during training, yielding superior performance when compared to previous methods including GNODE, GFINN, and SPNN discussed in Section 2. The metrics considered for this purpose will be mean absolute error (MAE) and mean squared error (MSE) defined in the usual way as the average $\ell^{1}$ (resp. squared $\ell^{2}$) error between the discrete states ${}x,\tilde{~{}x}\in\mathbb{R}^{n\times n_{s}}$. For brevity, many implementation details have been omitted here and can be found in Appendix B. An additional experiment showing the effectiveness of NMS in the presence of both full and partial state information can be found in Appendix D.
@@ 5.1 Two gas containers
where $(q,p)$ are the position and momentum of the separating wall, $S_{1},S_{2}$ are the entropies of the two subsystems, and the internal energies $E_{1},E_{2}$ are determined from the Sackur-Tetrode equation for ideal gases, $S_{i}/Nk_{B}=\ln\mathopen{}\mathclose{{}\left(\hat{c}V_{i}E_{i}^{3/2}}\right),1\leq i\leq 2.$ Here, $m$ denotes the mass of the wall, $2L$ is the total length of the system, and $V_{i}$ is the volume of the $i^{\mathrm{th}}$ container. As in Lee et al. (2021; 2022) $Nk_{B}=1$ and $\alpha=0.5$ fix the characteristic macroscopic unit of entropy while $\hat{c}=102.25$ ensures the argument of the logarithm defining $E_{i}$ is dimensionless. This leads to the total entropy $S(~{}x)=S_{1}+S_{2}$ and the total energy $E(~{}x)=(1/2m)p^{2}+E_{1}(~{}x)+E_{2}(~{}x),$ which are guaranteed to be nondecreasing and constant, respectively.
@@ 5.2 Thermoelastic double pendulum
Next, consider the thermoelastic double pendulum from Romero (2009) with 10-dimensional state variable ${}x=\begin{pmatrix}~{}q_{1}&~{}q_{2}&~{}p_{1}&~{}p_{2}&S_{1}&S_{2}\end{pmatrix}^{\intercal}$, which represents a highly challenging benchmark for metriplectic methods. The equations of motion in this case are given for $1\leq i\leq 2$ as
@@
The task in this case is prediction across initial conditions. As in Zhang et al. (2022), 100 trajectories are drawn from the ranges in Appendix B and integrated over the interval $[0,40]$ with $\Delta t=0.1$, with an 80/10/10 split for training/validation/testing. Here all compared models are trained using full state information. As seen in Table 1, NMS is again the most performant, although all models struggle to approximate the dynamics over the entire training interval. It is also notable that the training time of NMS is greatly decreased relative to GNODE and GFINN due to its improved quadratic scaling; a representative study to this effect is given in Appendix G.
@@ Appendix A Proof of Theoretical Results
This will be a direct consequence of Corollary 2.2 in Li (1996) provided we show that $\mathopen{}\mathclose{{}\left|~{}A}\right|\leq c\mathopen{}\mathclose{{}\left|~{}A}\right|_{\infty}$ for some $c>0$. To that end, if $\sigma_{1}\geq...\geq\sigma_{r}>0$ ($r\leq n$) denote the nonzero singular values of ${}A-\tilde{~{}A}$, it follows that for each ${}x\in K$,
@@
and therefore the desired inequality holds with $c=n$. Now, for any $\varepsilon>0$ it follows from Li (1996) that there exists a two layer network $\tilde{~{}A}$ with $m$ continuous derivatives such that $\sup_{~{}x\in K}\mathopen{}\mathclose{{}\left|~{}A-\tilde{~{}A}}\right|_{\infty}<\varepsilon/n$ and $\sup_{~{}x\in K}\mathopen{}\mathclose{{}\left|\nabla^{k}~{}A-\nabla^{k}\tilde{~{}A}}\right|_{\infty}<\varepsilon/n<\varepsilon$ for all $1\leq k\leq m$. Therefore, it follows that
@@
Now, by Corollary 2.2 in Li (1996), for any $\varepsilon>0$ there exists a two-layer neural network $\tilde{f}:K\to\mathbb{R}$ such that
@@ Appendix B Experimental and Implementation Details
This Appendix records additional details related to the numerical experiments in Section 5. For each benchmark problem, a set of trajectories is manufactured given initial conditions by simulating ODEs with known metriplectic structure. The metriplectic forms of these systems are given in the subsections below. For the experiments in Table 1, only the observable variables are used to construct datasets, since entropic information is assumed to be unknown. Algorithm 2 summarizes the training of the dynamics models used for comparison with NMS.
@@
For each compared method, integrating the ODEs is done via the Dormand–Prince method (dopri5) Dormand and Prince (1980) with relative tolerance $10^{-7}$ and absolute tolerance $10^{-9}$. The loss is evaluated by measuring the discrepancy between the ground truth observable states ${}x^{\text{o}}$ and the approximate observable states $\tilde{~{}x}^{\text{o}}$ in the mean absolute error (MAE) and mean squared error (MSE) metrics, which are defined in the usual way via the $\ell^{1},\ell^{2}$ norms respectively for true data ${}x(t_{i})$ and approximate data $\tilde{~{}x}(t_{i})$:
$MAE(\tilde{~{}x},~{}x)$ $=\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\mathopen{}\mathclose{{}\left|\tilde{~{}x}(t_{i})-~{}x(t_{i})}\right|_{1},\qquad MSE(\tilde{~{}x},~{}x)=\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\mathopen{}\mathclose{{}\left|\tilde{~{}x}(t_{i})-~{}x(t_{i})}\right|_{2}.$
The model parameters $\Theta$ (i.e., the weights and biases) are updated by using Adamax Kingma and Ba (2014) with an initial learning rate of 0.01. The number of training steps is set as 30,000, and the model parameters resulting in the best performance for the validation set are chosen for testing. Specific information related to the experiments in Section 5 is given in the subsections below.
@@ B.1 Two gas containers
The metriplectic equations of motion in the two gas containers example are given in terms of the state ${}x=\begin{pmatrix}q&p&S_{1}&S_{2}\end{pmatrix}^{\intercal}$. With the energy $E(~{}x)=(1/2m)p^{2}+E_{1}(~{}x)+E_{2}(~{}x)$ and entropy $S(~{}x)=S_{1}+S_{2}$ defined in Section 5, it follows that $\dot{~{}x}=~{}L\nabla E(~{}x)+~{}M(~{}x)\nabla S$ where $\nabla S=\begin{pmatrix}0&0&1&1\end{pmatrix}^{\intercal}$ and
$~{}L$ $=\begin{pmatrix}0&1&0&0\\ -1&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix},$ $~{}M(~{}x)=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&T_{1}^{-2}&-(T_{1}T_{2})^{-1}\\ 0&0&-(T_{1}T_{2})^{-1}&T_{2}^{-2}\end{pmatrix},$
$\nabla E(~{}x)$ $=\begin{pmatrix}-\mathopen{}\mathclose{{}\left(\frac{e^{2S_{1}}}{q^{5}}}\right)^{\frac{1}{3}}+\mathopen{}\mathclose{{}\left(\frac{e^{2S_{2}}}{(2-q)^{5}}}\right)^{\frac{1}{3}}\\ \frac{3}{2}p\\ T_{1}(~{}x)\\ T_{2}(~{}x)\end{pmatrix},$ $T_{i}(~{}x)=\partial_{S_{i}}E_{i}(~{}x).$
@@
In the experiment, a sequence of 100,000 timesteps is generated using the Runge–Kutta 4th-order (RK4) time integrator with a step size 0.001. The initial condition is given as ${}x=(1,2,103.2874,103.2874)$ following Shang and Öttinger (2020). The training/validation/test split is defined by $t_{\text{train}}=20$, $t_{\text{val}}=30$, and $t_{\text{test}}=100$. For a fair comparison, all considered models are set to have a similar number of model parameters, $\sim$2,000. The specifications of the network architectures are: